What is the Oxidation state of sulphur in thionic acid(h2s2o3)?Please explain it with the help of fundamental concept and with it's structure. 

Grade:12th pass

2 Answers

Arun
5 years ago
Dear Ashish
 
I am trying to give goosd explanation.
Na2S2O3
Oxidation state of 'S' you do not know, ok
Oxidation state for sodium = 1, So, Two sodium gets (1+1)= 2
Your question is oxidation state for sulphur, So unknown, There are two sulphur atoms are there, we will write as 2x
Oxidation state for oxygen = -2.There are there are three oxygen atoms, So
                                              (-2*3)= -6
Just apply in the avove molecular fomula,
Na2S2O3
(2+2x-6) = 0
(2-6)+2x= 0
-4+2x= 0
x " means sulphur atom = 2, So
Osidation state of sulphur is "2"
Ritik kumar
5 years ago
Oxidation state of sulpher is  +4 . Because S attached to other S with =bond lonely. and oxigen (with =bond),two OH gp attached with one of the S.so that one S has zero oxidn state and other has +4